Solve the following linear equations
(x - 5)/3 = (x - 3)/5 

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

(x-5)/3=(x-3)/5
⇒ 5(x-5) = 3(x-3)
⇒ 5x-25 = 3x-9
⇒5x-3x=-9+25
⇒2x=16
⇒x=8

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
(X-5)/3 = (X-3)/5
By cross multiplication,
5(X-5) = 3(X-3)
5X-25 = 3X-9
2X = 16
X = 8
 
Hope it helps! 
Thank you ✌️